1. Rocket observations show that dust particles in Earth's upper atmosphere are often electrically charged.

a) Find the distance separating two dust particles (in m) if each has a charge of +e and the Coulomb force between them has magnitude 1.30 ✕ 10−14 N.

b)Calculate the mass (in kg) of one of the dust particles if this Coulomb force would accelerate it at 2.60 ✕ 108 m/s2. (In the upper atmosphere, effects from other nearby charges typically result in a small net force and acceleration.)

2. Two small identical conducting spheres are placed with their centers 0.36 m apart. One is given a charge of 12 ✕ 10−9 C, the other a charge of −23 ✕ 10−9 C.
(a) Find the electrostatic force exerted on one sphere by the other.

b)The spheres are connected by a conducting wire. Find the electrostatic force between the two after equilibrium is reached, where both spheres have the same charge.
Answer #1

1.

a)

F = k q^2 / r^2

1.3* 10^-14 = 9* 10^9* (1.6* 10^-19)^2 / r^2

r = 1.33* 10^-7 m

b)

ma = F

m * 2.6* 10^8 = 1.3* 10^-14

m = 5* 10^-23 kg

======

plz post other problem separately

Comment before rate in case any doubt, will reply for sure.. goodluck